AP4 ) How much rust (Fe 2 O 3 ) will be formed, when 100 kg of iron have completely rusted away? ( Atomic weights: Fe = 56 ; O = 16; H = 1)

Grade:12th pass

1 Answers

Naveen Kumar
9 years ago
The reaction can be given as:
4Fe..+.......3O2.............>2Fe2O3
so 100kg iron means.....100*1000g Fe
....................................(100000/56) mol Fe
From reaction we can see that 4 mol Fe will give 2 mol Fe2O3
So,.........................................1mol.................2/4mol..........
So,number of mol of Fe2O3 formed=(2/4)*(100000/56)=m(say)
Molecular weight of Fe2O3=(56*2)+(16*3)=160g/mol
Hence mass of Fe2O3 formed=m*160g